# Hyperparameter Transfer with Mixture-of-Experts Layers

### v1

@@ 5 Empirical results and findings
Figure 5: Parts (1) and (2): At a fixed token budget of 1B, scaling up from the base model (1 activated expert and expert $\alpha_{\operatorname{ffn}}=1$), increasing the number of experts is more parameter-efficient than expert size. Error bars are (min, median, max) out of 4 seeds. Part (3): At a 5B token horizon with GPT2-small activated and cosine LR cool-down to zero, having more activated experts (and thus smaller experts) monotonically improves performance. Error bars for $n_{\operatorname{act}}\leq 4$ are (min, median, max) out of 3 seeds.
Our experiments focus on MoE architectures described in Section˜3.1, and we defer training details to Appendix A.
We consider two different expert activation sparsity on two different natural language datasets: (i) Sparsity $\kappa=1/4$, on the FineWeb dataset (Penedo et al., 2024); and (ii) Sparsity $\kappa=1/12$, on the C4 (Colossal Clean Crawled Corpus) dataset (Raffel et al., 2020).
In all of our experiments under a fixed token budget, we use a total of (roughly) 1B tokens divided into 2000 batches of 500K tokens and context length 1024. We use an LR scheduler with a linear warmup phase for the first 1000 steps and stable (constant) LR for the latter half. While a typical LR schedule does not warm up for half of the training iterations, our goal with the fixed token budget experiments is to model an early checkpoint or a larger run, which often contains 0.5B (or more) tokens during the warm-up phase.
For each parameter group, we also tune constant-scale multipliers on the learning rate and initialization. Without tuning constant multipliers, we found that (1) nontrivial performance was left on the table and (2) training dynamics (e.g. load balancing loss) can become unstable even at around the optimal HP. See Appendix D.1 for details
@@ 5.1 Experiment findings for fixed token budget
Finding 1.2. On the upscaled optimal HPs, loss profiles in early iterations collapse to that of the base model as width, number of experts, and expert sizes increase.
@@ 5.2 Experiments on larger token-horizon
Finding 2.1. The optimal HPs found from small models enable stable training on longer total token horizons and achieve competitive loss (against active parameter-matched dense models).
Fixing the architecture of the total activated model to match that of the Nano-GPT implementation of GPT2-small (124M) (and having more total parameters), we report a competitive loss curve via zero-shot HPs (Figure˜1) compared to checkpoints of the (dense) GPT2-Fineweb speedrun (Jordan & contributors, 2025) under the same batch size. See also Figure˜16 for running 7.5B tokens on GPT2-medium (355M activated). For training with a large number of steps, while fixing a stable LR after warmup still yields stable and converging training, including a LR decay (even the simplest cosine decay to zero) empirically improves final validation performance in our experiments.
@@ 5.3 Tradeoff between expert count versus expert size
We can now study architectural properties such as the tradeoff between expert count and expert size while fixing sparsity and number of parameters.
Finding 3.1. An increasing number of experts while fixing the total parameter improves the final model performance.
We justify it via Figure˜5 and in Apdx D.3 for smaller models trained on a longer token horizon. While previous works reported benefits of more smaller experts (Krajewski et al., 2024; Liu et al., 2024; Boix-Adsera & Rigollet, 2025), our HP transfer enables such a comparison to be fair and cheap.
@@ 6 Conclusion and future directions
    As training dynamics in early stages (our 1B scale experiments) versus later stages can be very different, it would be interesting to expand the HP transfer to longer training horizons that are practical.
@@
    We’d also like to remark that it is not clear whether “compute-optimal” token horizon rules for dense models are practical for MoEs (Clark et al., 2022), as MoEs (a) achieve significantly better performance compared to dense models under the same FLOP budget, so exact conclusions of Chinchilla may not apply and (b) even fixing FLOP, take up significantly more demanding hardware resources (time and memory) due to sparsity.
    We take our work on HP transfer as a first step to derive scaling law arguments for MoE models in general, at least at a fixed token horizon (see (Clark et al., 2022; Krajewski et al., 2024; Li et al., 2025) for prior literature on fitted MoE scaling laws in different settings).
@@
    Finally, we left out discussions of other hyperparameters such as batch size, Adam betas, and LR scheduling, all of which are often heavily interconnected and can have a more significant impact as data scale expands.
@@ Appendix A More experiment details
Our experiments are run on H100s and bit-wise deterministic. See supplementary materials for implementation details.
Appendix B Learning rate for biases
@@
    print(len(mask_top_a & mask_top_b) / active)
@@ Appendix C Weight decay
Figure 7: Weight decay with independent scaling preserving $\eta\lambda$ (x-axis here is the re-scaled $\tau_{\operatorname{EMA}}$). While we do see preliminary transfer behaviors across model scales when using independent weight decay, we do not observe significant benefit from a nontrivial WD.
While weight decay (WD) is an important hyperparameter in AdamW (Loshchilov & Hutter, 2019), the exact empirical effect in the context of HP transfer is not well understood. For instance, (Kosson et al., 2025) found that independent weight decay, keeping $\eta\cdot\lambda$ constant (across model scales) such that in every step a constant fraction of weights gets whitened, enables HP transfer. On top of that, (Dey et al., 2025) argued that this independent weight decay $\eta\lambda$ may require depth adjustments according to the depth scaling exponent. However, another line of empirical works (Wang & Aitchison, 2025; Bergsma et al., 2025a) suggests that the scaled invariant should be the effective EMA timescale $\tau=T\eta\lambda$ where $T$ is the number of steps trained. This is contradictory with independent weight decay when $T$ is connected to model parameters, such as width or depth (often the practical case for “compute-optimality” (Kaplan et al., 2020)). Finally, works such as (Defazio, 2025) suggested WD scheduling, another variable to be considered.
In our experiments with weight decay in a limited scope, we found that at the 1B token horizon with $T=2000$ steps, a nontrivial weight decay does not significantly outperform having zero weight decay altogether, and hence all our experiments in the main text are done without weight decay. See Figure˜7.
@@ D.1 Constant-scale HP tuning for the base model
In (Yang & Hu, 2022; Yang et al., 2022), the authors pointed out that good HP transfer is crucially based on the “close to optimality” of the scaling. Without specific care, HPs found from tuning small models can be “contaminated” by finite-width effects, which fail to yield useful transfer (Figure 17, (Ghosh et al., 2025)). This motivates tuning constant-scaled multipliers on relevant HPs, whose benefits for dense models were remarked in (Everett et al., 2024; Mlodozeniec et al., 2025; Ghosh et al., 2025) that not only promote better loss (in the base model) but likely contribute to more reliable transfer. In our experiments with MoE layers, we found that balancing constant-scaled hyperparameters not only leads to better performance, but it is specifically crucial for stable training dynamics (across different random seeds) on large learning rates. In practice, we enable two constant multipliers for each MoE parameter groups for learning rate and initialization. While each group of weights has a three relevant HPs: learning rate, initialization, and multiplier, in the so-called abc-parameterization (Yang & Hu, 2022), only two degrees of freedom are present. We will take all weight multipliers to be one, so only initializations and learning rates are at play. We report loss results on ablation in Figure˜9 and load balancing ablation in Figure˜9.
@@ D.3 More experiments
    In Figure˜16, we compare our results ran on GPT2-medium (355M). While recorded (Jordan & contributors, 2025) Speedrun logs for this architecture only exist after the 124M base variant was highly optimized (which vastly improves Figure˜1), we still manage to outperform the (tuned) llm.c baseline (Karpathy, 2023) easily.
@@ Appendix E Dynamical mean field theory
In this section, we derive the dynamical mean field theory equations for the large width, large expert count, and large depth limit of this model. We assume a deep residual network consisting of $L$ MoE layers without attention. We further focus on gradient flow in the present analysis but this can be easily relaxed to discrete time SGD (Bordelon & Pehlevan, 2022; Bordelon et al., 2023).
@@
$\mathcal{Z}_{\text{res}}=\int\prod_{\ell}$ $\mathcal{D}\bar{\chi}^{\ell}\mathcal{D}\widehat{\bar{\chi}}^{\ell}\mathcal{D}\bar{\xi}^{\ell}\mathcal{D}\widehat{\bar{\chi}}^{\ell}\mathcal{D}h^{\ell}\mathcal{D}\widehat{h}^{\ell}\mathcal{D}g^{\ell}\mathcal{D}\widehat{g}^{\ell}\exp\left(-\frac{\alpha_{\star}}{2}\sum_{\ell}\left[\widehat{\bar{\chi}}^{\ell}\widehat{\bar{\chi}}^{\ell}M^{\ell}_{\sigma\sigma C_{\phi}}+\widehat{\bar{\xi}}^{\ell}\widehat{\bar{\xi}}^{\ell}M^{\ell}_{\sigma\sigma C_{g}}\right]\right)$
@@
$\alpha_{\star}=\frac{N}{N_{e}E}$ (29)
@@
This is satisfied for many common scaling strategies. For instance, if FFN ratio $N_{e}/N$ is fixed and $E$ also diverges (at any rate) then this condition is satisfied as $\alpha_{\star}=0$. Similarly, if $E/N$ is fixed and $N_{e}$ diverges (at any rate) then $\alpha_{\star}=0$.
@@
The remaining saddle point equations give $\widehat{C}=0$ and $\widehat{M}=0$. The response functions can be rearranged as derivatives through integration by parts (Crisanti & Sompolinsky, 2018; Mignacco et al., 2020; Bordelon & Pehlevan, 2022, 2026)
@@
The DMFT single site equations are the following. For the residual stream (reminder that $\alpha_{\star}=\lim_{N\to\infty}\frac{N}{N_{e}(N)E(N)}$ is bounded).
@@
$u^{\ell}(\bm{x},t)\sim\mathcal{GP}\left(0,\alpha_{\star}M^{\ell}_{\sigma\sigma C_{\phi_{k}}}(\bm{x},\bm{x}^{\prime},t,t^{\prime})\right)\ ,\ r^{\ell}(\bm{x},t)\sim\mathcal{GP}\left(0,\alpha_{\star}M^{\ell}_{\sigma\sigma C_{g_{k}}}(\bm{x},\bm{x}^{\prime},t,t^{\prime})\right)$ (47)
@@
Next, for the expert distribution, we have
$p_{k}^{\ell}(\bm{x},t)=\gamma_{0}\int dt^{\prime}\mathcal{A}_{k}^{\ell}(\bm{x}^{\prime},t^{\prime})\dot{\sigma}_{k}^{\ell}(\bm{x}^{\prime},t^{\prime})H^{\ell}(\bm{x},\bm{x}^{\prime},t,t^{\prime})$
$b_{k}(t)=b_{k}(0)+\gamma_{0}\int dt^{\prime}\left(\kappa-\mathbb{E}_{\bm{x}}1_{q^{\ell}(\bm{x},t)\geq q^{\ell}_{\star}(\bm{x},t)}\right)$
$\mathcal{A}_{k}^{\ell}(\bm{x},t)=C^{\ell}_{\phi_{k}\xi_{k}}(\bm{x},\bm{x},t,t)+\gamma_{0}\mathbb{E}_{\bm{x}^{\prime}}\int dt^{\prime}\sigma_{k}^{\ell}(\bm{x}^{\prime},t^{\prime})C_{g}^{\ell+1}(\bm{x},\bm{x}^{\prime},t,t^{\prime})C^{\ell}_{\phi_{k}}(\bm{x},\bm{x}^{\prime},t,t^{\prime})$
$\sigma_{k}^{\ell}(\bm{x},t)=1_{q^{\ell}(\bm{x},t)\geq q^{\ell}_{\star}(\bm{x},t)}\sigma(p_{k}^{\ell}(\bm{x},t))$ (48)
These equations define the averaging operation over experts $\left[\right]$. The main source of disorder from the initial condition arises from the random initial biases $b_{k}(0)$ for the router. The dynamics of neurons within expert have the form
@@
Similarly, we have an SDE description for the residual stream variables
@@
$\left<du(\tau,\bm{x},t)du(\tau^{\prime},\bm{x}^{\prime},t^{\prime})\right>=d\tau d\tau^{\prime}\delta(\tau-\tau^{\prime})M_{\sigma\sigma C_{\phi}}(\tau,\bm{x},\bm{x}^{\prime},t,t^{\prime})$ (52)

### v2

@@ 5 Empirical results and findings
Figure 5: Parts (1) and (2): At a fixed token budget of 1B, when scaling up from the base model ($n_{\operatorname{act}}=\alpha_{\operatorname{ffn}}=1$), increasing the number of experts is more parameter-efficient than expert size. Error bars are (min, median, max) out of 4 seeds. Part (3): At a 5B token horizon with GPT2-small activated and cosine LR cool-down to zero, having more activated experts (and thus inverse proportionally smaller experts) monotonically improves performance. Error bars for $n_{\operatorname{act}}\leq 4$ are (min, median, max) out of 3 seeds.
Our experiments focus on MoE architectures described in Section˜3.1, and we defer training details to Appendix A. We consider two different sparsities $\kappa$ on two different natural language datasets: (i) $\kappa=1/4$ on the FineWeb dataset (Penedo et al., 2024); and (ii) $\kappa=1/12$ on the C4 (Colossal Clean Crawled Corpus) dataset (Raffel et al., 2020).
In all of our experiments with a fixed token budget, we use a total of (roughly) 1B tokens divided into 2000 batches of 500K tokens and context length 1024. We use an LR scheduler with a linear warmup phase for the first 1000 steps and stable (constant) LR for the latter half. While a typical LR schedule does not warm up for half of the training iterations, our goal with the fixed token budget experiments is to model an early checkpoint or a larger run, which often contains 0.5B (or more) tokens during the warm-up phase.
For each parameter group, we also tune constant-scale multipliers on the learning rate and initialization. Without tuning constant multipliers, we found that (1) nontrivial performance was left on the table and (2) training dynamics (e.g. load balancing loss) can become unstable even at around the optimal HP. See Appendix D.1 for details.
@@ 5.1 Experiment findings for fixed token budget
Finding 1.2. On the upscaled optimal HPs, loss profiles in early iterations collapse to those of the base model as width, number of experts, and expert sizes increase.
@@ 5.2 Experiments on larger token-horizon
Finding 2.1. Optimal HPs found in small models enable stable training on longer token horizons and achieve competitive loss (against active parameter-matched dense models).
Fixing the architecture of the total activated model to match that of the Nano-GPT implementation of GPT2-small (124M) (and having more total parameters), we report a competitive loss curve via zero-shot HPs (Figure˜1) compared to checkpoints of the (dense) GPT2-Fineweb speedrun (Jordan and contributors, 2025) under the same batch size. See also Figure˜16 for running 7.5B tokens on GPT2-medium (355M activated). For training with a large number of steps, while fixing a stable LR after warmup still yields stable and converging training, including a LR decay (even the simplest cosine decay to zero) empirically improves final validation performance in our experiments.
@@ 5.3 Tradeoff between expert count versus expert size
We can now study architectural choices such as the tradeoff between expert count and expert size while fixing sparsity and number of (total and active) parameters.
Finding 3.1. Increasing number of experts at fixed parameter count improves final model performance.
We justify this via Figure˜5 (for short horizon training runs) and further in Apdx D.3 (small models on long horizons). While previous works have reported benefits of more smaller experts (Krajewski et al., 2024; Liu et al., 2024; Boix-Adsera and Rigollet, 2025), our HP transfer results enables such architectural comparison results to be fair (across dimensions) and cheap (without needing to sweep HPs).
@@ 6 Conclusion and future directions
    As training dynamics in early stages (such as our experiments in 1B token scale) versus later stages (e.g., the “compute-optimal” horizon) can be very different, it would be interesting to expand the HP transfer to longer training horizons that are practical.
@@
    We’d also like to remark that it is not known what “compute-optimal” rules can be practically applied for MoEs (Clark et al., 2022), as (a) Mixture-of-Expert layers achieve significantly better performance compared to dense models under the same FLOP budget, so Chinchilla exponents may not apply, and (b) even when FLOP-matched, MoEs take up significantly more demanding hardware resources due to sparsity and routing, suggesting the necessity of a better way of evaluating compute.
    We take our present work on HP transfer as a first step to derive practical scaling laws for MoE models (see (Clark et al., 2022; Krajewski et al., 2024; Li et al., 2025; Zhao et al., 2025) for prior MoE laws in different settings).
@@
    Finally, we left out discussions of hyperparameters beyond learning rate and init standard deviation, such as batch size, Adam betas, and LR schedules, all of which are often interconnected and can have a more significant impact as data scale expands.
Acknowledgements
T.J. is supported by DARPA AIQ-HR001124S0029. B.B. thanks the the Center of Mathematical Sciences and Applications (CMSA) of Harvard University for support. C.P. is supported by an NSF CAREER Award (IIS-2239780), DARPA grants DIAL-FP-038 and AIQ-HR00112520041, the Simons Collaboration on the Physics of Learning and Neural Computation, and the William F. Milton Fund from Harvard University. This work has been made possible in part by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence. B.H. is grateful for support from a 2024 Sloan Fellowship in Mathematics, NSF CAREER grant DMS-2143754, and NSF grant DMS-2133806, and DARPA AIQ-HR001124S0029. We thank Mithril for providing compute resources.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A More experiment details
Our experiments are run on H100s and are bit-wise deterministic. Our code is available at https://github.com/Petyrrrrr/MuP_MOE/tree/complete.
Appendix B Heuristic LR derivation for expert biases
@@
    print(len(mask_top_a & mask_top_b) / active) #symmetric difference portion
@@ Appendix C Weight decay
Figure 7: Weight decay with independent scaling preserving $\eta\lambda$ (x-axis is the re-scaled $\tau_{\operatorname{EMA}}$). While we see preliminary transfer behaviors across scales when using independent weight decay, we do not observe significant benefit from a nontrivial WD in our scale.
While weight decay (WD) is an important hyperparameter in AdamW (Loshchilov and Hutter, 2019), the exact empirical effect in the context of HP transfer is not well understood. For instance, (Kosson et al., 2025) found that independent weight decay, keeping $\eta\cdot\lambda$ constant (across model scales) such that in every step a constant fraction of weights gets whitened, enables HP transfer. On top of that, (Dey et al., 2025) argued that this independent weight decay $\eta\lambda$ may require depth adjustments according to the depth scaling exponent. However, another line of empirical works (Wang and Aitchison, 2025; Bergsma et al., 2025a) suggests that the scaled invariant should be the effective EMA timescale $\tau=T\eta\lambda$ where $T$ is the number of steps trained. This is contradictory with independent weight decay when $T$ is connected to model parameters, such as width or depth (often the practical case for “compute-optimality” (Kaplan et al., 2020)). Finally, works such as (Defazio, 2025) suggested WD scheduling, another variable to be considered.
In our experiments with weight decay in a limited scope, we found that at the 1B token horizon with $T=2000$ steps, a nontrivial weight decay does not significantly outperform having zero weight decay in our horizon, and hence all our experiments in the main text are done without weight decay. See Figure˜7.
@@ D.1 Constant-scale HP tuning for the base model
In (Yang and Hu, 2022; Yang et al., 2022), the authors pointed out that good HP transfer is crucially based on the “close to optimality” of the scaling. Without specific care, HPs found from tuning small models can be “contaminated” by finite-width effects, which fail to yield useful transfer (Figure 17, (Ghosh et al., 2025)). This motivates tuning constant-scaled multipliers on relevant HPs, whose benefits for dense models were remarked in (Everett et al., 2024; Mlodozeniec et al., 2025; Ghosh et al., 2025) that not only promote better loss (in the base model) but likely contribute to more reliable transfer. In our experiments with MoE layers, we found that balancing constant-scaled hyperparameters not only leads to better performance, but it is specifically crucial for stable training dynamics (across different random seeds) on large learning rates. In practice, we enable two constant multipliers for each MoE parameter groups for learning rate and initialization. While each group of weights has a three relevant HPs: learning rate, initialization, and multiplier, in the so-called abc-parameterization (Yang and Hu, 2022), only two degrees of freedom are present. We will take all weight multipliers to be one, so only initializations and learning rates are at play. We report loss results on ablation in Figure˜9 and load balancing ablation in Figure˜9.
@@ D.3 More experiments
    In Figure˜16, we compare our results ran on GPT2-medium (355M). While recorded (Jordan and contributors, 2025) Speedrun logs for this architecture only exist after the 124M base variant was highly optimized (which vastly improves Figure˜1), we still manage to outperform the (tuned) llm.c baseline (Karpathy, 2023) easily.
@@ Appendix E Dynamical mean field theory
In this section, we derive the dynamical mean field theory equations for the large width, large expert count, and large depth limit of this model. We assume a deep residual network consisting of $L$ MoE layers without attention. We further focus on gradient flow in the present analysis but this can be easily relaxed to discrete time SGD (Bordelon and Pehlevan, 2022; Bordelon et al., 2023).
@@
$\mathcal{Z}_{\text{res}}=\int\prod_{\ell}$ $\mathcal{D}\bar{\chi}^{\ell}\mathcal{D}\widehat{\bar{\chi}}^{\ell}\mathcal{D}\bar{\xi}^{\ell}\mathcal{D}\widehat{\bar{\chi}}^{\ell}\mathcal{D}h^{\ell}\mathcal{D}\widehat{h}^{\ell}\mathcal{D}g^{\ell}\mathcal{D}\widehat{g}^{\ell}\exp\left(-\frac{\alpha_{\star}L}{2}\sum_{\ell}\left[\widehat{\bar{\chi}}^{\ell}\widehat{\bar{\chi}}^{\ell}M^{\ell}_{\sigma\sigma C_{\phi}}+\widehat{\bar{\xi}}^{\ell}\widehat{\bar{\xi}}^{\ell}M^{\ell}_{\sigma\sigma C_{g}}\right]\right)$
@@
$\alpha_{\star}\equiv\frac{N}{N_{e}EL}$ (29)
@@
This is satisfied for many common scaling strategies. For instance, if FFN ratio $N_{e}/N$ is fixed and $E$ also diverges (at any rate) then this condition is satisfied as $\alpha_{\star}=0$. Similarly, if $E/N$ is fixed and $N_{e}$ diverges (at any rate) then $\alpha_{\star}=0$. We consider simultaneously diverging depth below.
@@
The remaining saddle point equations give $\widehat{C}=0$ and $\widehat{M}=0$. The response functions can be rearranged as derivatives through integration by parts (Crisanti and Sompolinsky, 2018; Mignacco et al., 2020; Bordelon and Pehlevan, 2022, 2026)
@@
The DMFT single site equations are the following. For the residual stream (reminder that $\alpha_{\star}=\lim_{N\to\infty}\frac{N}{N_{e}(N)E(N)L}$ is bounded).
@@
$u^{\ell}(\bm{x},t)\sim\mathcal{GP}\left(0,\alpha_{\star}LM^{\ell}_{\sigma\sigma C_{\phi_{k}}}(\bm{x},\bm{x}^{\prime},t,t^{\prime})\right)\ ,\ r^{\ell}(\bm{x},t)\sim\mathcal{GP}\left(0,\alpha_{\star}LM^{\ell}_{\sigma\sigma C_{g_{k}}}(\bm{x},\bm{x}^{\prime},t,t^{\prime})\right)$ (47)
@@
Next, for the expert distribution, we have the following DMFT equations for router preactivation $p$
$p^{\ell}(\bm{x},t)=\gamma_{0}\int dt^{\prime}\mathcal{A}^{\ell}(\bm{x}^{\prime},t^{\prime})\dot{\sigma}^{\ell}(\bm{x}^{\prime},t^{\prime})H^{\ell}(\bm{x},\bm{x}^{\prime},t,t^{\prime})$
$b(t)=b(0)+\gamma_{0}\int dt^{\prime}\left(\kappa-\mathbb{E}_{\bm{x}}1_{q^{\ell}(\bm{x},t)\geq q^{\ell}_{\star}(\bm{x},t)}\right)$
$\mathcal{A}^{\ell}(\bm{x},t)=C^{\ell}_{\phi_{k}\xi_{k}}(\bm{x},\bm{x},t,t)+\gamma_{0}\mathbb{E}_{\bm{x}^{\prime}}\int dt^{\prime}\sigma^{\ell}(\bm{x}^{\prime},t^{\prime})C_{g}^{\ell+1}(\bm{x},\bm{x}^{\prime},t,t^{\prime})C^{\ell}_{\phi}(\bm{x},\bm{x}^{\prime},t,t^{\prime})$
$\sigma^{\ell}(\bm{x},t)=1_{q^{\ell}(\bm{x},t)\geq q^{\ell}_{\star}(\bm{x},t)}\ \sigma(p^{\ell}(\bm{x},t))$ (48)
These equations define the averaging operation over experts $\left[\right]$. The main source of disorder from the initial condition arises from the random initial biases $b(0)$ for the router. Lastly the mean field dynamics of each neuron within the experts have the form
@@
Similarly, we have an SDE description for the residual stream variables which are driven by an additional Brownian motion $du(\tau,\bm{x},t)$
@@
$\left<du(\tau,\bm{x},t)du(\tau^{\prime},\bm{x}^{\prime},t^{\prime})\right>=d\tau d\tau^{\prime}\delta(\tau-\tau^{\prime})M_{\sigma\sigma C_{\phi}}(\tau,\bm{x},\bm{x}^{\prime},t,t^{\prime}).$ (52)